OpenStudy (anonymous):

What are the odds of picking an 'm' in the word Alabama? Isn't the answer 1:6? The probability would be 1/7, correct? If you google this question most people are saying the odds are 1:7, but I don't think thats correct...am I right?

OpenStudy (bibby):

there are 7 letters in alabama, and so the probability would be \(\frac{1}{7}\)

OpenStudy (anonymous):

I'm asking for the odds

OpenStudy (igreen):

There are 7 total letters in the word Alabama. There is 1 'm', in alabama, therefore the probability is \(\dfrac{1}{7}\)

OpenStudy (anonymous):

Wouldn't the odds be 1:6

OpenStudy (igreen):

Probability is the same thing as 'odds'.

OpenStudy (igreen):

No, it's not 1:6, because there are SEVEN total letters and NOT SIX.

OpenStudy (igreen):

So it's 1:7 or \(\dfrac{1}{7}\).

OpenStudy (anonymous):

but @iGreen It even says right in my unit Odds are # of desirable outcomes/ # of undesirable outcomes and that probability is # of desirable outcomes/# of possible outcomes

OpenStudy (anonymous):

so with following this wouldnt the probability be 1/7 and the odds be 1:6?

OpenStudy (igreen):

It's incorrect, the probability is 1/7.

OpenStudy (igreen):

@Duffy0121

OpenStudy (anonymous):

I think I should go off of what my unit says, Probability and odds are different, let me pull you an example.

OpenStudy (anonymous):

Odds is defined as the ratio of the number of ways a successful event can occur to the number of ways the event can’t occur.

OpenStudy (anonymous):

Suppose you have a six sided die (singular of dice). What are the odds and probability of rolling a number greater than or equal to 5? Odds is defined as success : failures. There are two numbers greater than or equal to 5. Those numbers are 5 and 6. The failures would be the numbers that are not greater than or equal to 5. Those numbers are 1, 2, 3, and 4. This means that odds is 2 : 4 or 1 : 2.

OpenStudy (anonymous):

So see what I mean @iGreen

OpenStudy (igreen):

Yes, I do..if odds are Success : Failures, then 1 : 6 does seem correct.